### v2

# Vogelpothsweg 87, D-44227 Dortmund, Germany

11institutetext: Prof. Dmitri Kuzmin 22institutetext: Institute of Applied Mathematics (LS III), TU Dortmund University
Vogelpothsweg 87, D-44227 Dortmund, Germany
33institutetext: 33email: kuzmin@math.tu-dortmund.de
Convex limiting for finite elements and its relationship to residual distribution
Dmitri Kuzmin
MSC2020: 62M12; 65M60.
1 Introduction
The design of a finite element scheme that ensures preservation of local and/or global bounds for numerical solutions to a hyperbolic problem typically involves construction and local adjustment of a dissipative shock capturing term. In most cases, a finite element counterpart of the local Lax–Friedrichs (Rusanov) method is corrected by adding limited antidiffusive terms. In edge-based schemes, the additional terms consist of fluxes between pairs of neighboring nodes guermond2018 ; convex ; fctbook ; kuzmin2010 ; luo1994 . In element-based alternatives, diffusive and antidiffusive correction terms are assembled from element vectors that possess the zero sum property dobrev2018 ; preprint ; kuzmin2020 ; lohmann2017 . A comprehensive review of the state of the art can be found in the recent book book .
The element-based approach traces its origins to the flux-corrected transport method proposed by Löhner et al. lohner1987 ; lohner1988 . Moreover, it is closely related to residual distribution (RD) schemes that achieve the algebraic stabilization effect by manipulating element contributions in a locally conservative manner abgrall2006 ; abgrall2017 ; abgrall2019 ; abgrall2010 . In this note, we review the recent progress in the development of element-based convex limiting / RD tools for continuous Galerkin methods using linear finite elements.
2 Element-based Rusanov method
Let $u(\mathbf{x},t)$ denote an exact solution of the hyperbolic conservation law or system
$\frac{\partial u}{\partial t}+\nabla\cdot\mathbf{f}(u)=0\quad\mbox{in}\ \Omega\times(0,T].$ (1)
We assume that periodic boundary conditions are imposed on $\partial\Omega$ and that there exists a convex invariant set $\mathcal{G}$ such that $u\in\mathcal{G}$ a.e. in $\Omega\times(0,T]$. In the context of gas dynamics, all states belonging to $\mathcal{G}$ have positive density and internal energy.
A continuous piecewise-linear ($\mathbb{P}_{1}$) finite element approximation $u_{h}=\sum_{j=1}^{N_{h}}u_{j}\varphi_{j}$ to $u$ is called invariant domain preserving (IDP) or positivity preserving if $u_{j}\in\mathcal{G}$ for $j=1,\dots,N_{h}$. The Lagrange basis functions $\varphi_{i}$ are associated with vertices $\mathbf{x}_{i}$ of a conforming mesh $\mathcal{T}_{h}$. The elements (cells) of that mesh are denoted by $E_{1},\ldots,E_{h}$. The global indices of nodal points $\mathbf{x}_{i}$ belonging to $K_{e}$ are stored in the set $\mathcal{N}^{e}$. The indices of elements $K_{e}$ containing $\mathbf{x}_{i}$ are stored in the set $\mathcal{E}_{i}$.
The standard Galerkin discretization of the hyperbolic problem (1) is given by
$\sum_{e\in\mathcal{E}_{i}}m_{ij}^{e}\frac{\mathrm{d}u_{j}}{\mathrm{d}t}=\sum_{e\in\mathcal{E}_{i}}\int_{K_{e}}\nabla\varphi_{i}\cdot\mathbf{f}(u_{h})\,\mathrm{d}\mathbf{x},\qquad i=1,\ldots,N_{h},$ (2)
where $m_{ij}^{e}=\int_{K_{e}}\varphi_{i}\varphi_{j}\,\mathrm{d}\mathbf{x}$ is the contribution of $K_{e}$ to the entry $m_{ij}=\sum_{e\in\mathcal{E}_{i}\cap\mathcal{E}_{j}}m_{ij}^{e}$ of the consistent mass matrix $M_{C}=\{m_{ij}\}_{i,j=1}^{N_{h}}$.
To construct a low-order IDP approximation to (2), we lump the mass matrix, use the midpoint rule for the volume integrals on the right-hand side, and add Rusanov artificial viscosity of the form employed in abgrall2006 . These manipulations lead to
$m_{i}\frac{\mathrm{d}u_{i}}{\mathrm{d}t}=\sum_{e\in\mathcal{E}_{i}}[d^{e}(\bar{u}^{e}-u_{i})-\mathbf{f}(\bar{u}^{e})\cdot\mathbf{c}_{i}^{e}],\qquad i=1,\ldots,N_{h},$ (3)
where $\bar{u}_{e}$ is the arithmetic mean of $u_{i},\ i\in\mathcal{N}^{e}$ and $\mathbf{c}_{i}^{e}=-\int_{K_{e}}\nabla\varphi_{i}\,\mathrm{d}\mathbf{x}$. The Rusanov viscosity coefficient $d^{e}=\max_{i\in\mathcal{N}^{e}}\lambda_{i}^{e}|\mathbf{c}_{i}^{e}|$ is defined using the maximum speeds $\lambda_{i}^{e}$ of Riemann problems with flux functions $\mathbf{f}(u)\cdot\mathbf{c}_{i}^{e}/|\mathbf{c}_{i}^{e}|$ and initial states $(\bar{u}^{e},u_{i})$.
The element-based low-order method (3) is a slightly modified version of the one derived in (book, , Sec. 4.6.1) using a trapezoidal rule approximation. If (3) is discretized in time using an explicit strong stability preserving (SSP) Runge–Kutta method, then each low-order forward Euler stage can be written as
$u_{i}^{L}=\Big(1-\frac{2\Delta t}{m_{i}}\sum_{e\in\mathcal{E}_{i}}d^{e}\Big)u_{i}+\frac{2\Delta t}{m_{i}}\sum_{e\in\mathcal{E}_{i}}d^{e}\bar{u}_{i}^{e}.$ (4)
Similarly to the bar states of the edge-based version analyzed in guermond2016 , the state
$\bar{u}_{i}^{e}=\frac{\bar{u}^{e}+u_{i}}{2}-\frac{(\mathbf{f}(\bar{u}^{e})-\mathbf{f}(u_{i}))\cdot\mathbf{c}_{i}^{e}}{2d_{i}^{e}},\qquad e\in\mathcal{E}_{i}$
represents a spatially averaged exact solution of the associated Riemann problem. Thus, $\bar{u}^{e},u_{i}\in\mathcal{G}$ implies $\bar{u}_{i}^{e}\in\mathcal{G}$ for any convex invariant set $\mathcal{G}$. If the time step $\Delta t$ satisfies the CFL-like condition $\frac{2\Delta t}{m_{i}}\sum_{e\in\mathcal{E}_{i}}d^{e}\leq 1$, then $u_{i}^{L}$ is a convex combination of the IDP states $u_{i}$ and $\bar{u}_{i}^{e},\ e\in\mathcal{E}_{i}$. Hence, the fully discrete scheme is IDP.
3 Residual distribution
To interpret (3) as a low-order RD method derived from (2), we notice that both semi-discrete schemes can be written in the generic form (cf. (book, , Chapter 4))
$m_{i}\frac{\mathrm{d}u_{i}}{\mathrm{d}t}=\sum_{e\in\mathcal{E}_{i}}r_{i}^{e},$ (5)
where $r_{i}^{e}$ denotes the contribution of $K_{e}$ to the steady-state residual. Using integration by parts, we deduce that (2) is equivalent to (5) for $r_{i}^{e}=r_{i}^{e,H}$ given by
$r_{i}^{e,H}=\int_{K_{e}}\varphi_{i}(\dot{u}_{i}-\dot{u}_{h})\,\mathrm{d}\mathbf{x}-\int_{K_{e}}\varphi_{i}\nabla\cdot\mathbf{f}(u_{h})\,\mathrm{d}\mathbf{x}.$
The coefficients $\dot{u}_{i}=\frac{\mathrm{d}u_{i}}{\mathrm{d}t}$ of $\dot{u}_{h}=\frac{\mathrm{d}u_{h}}{\mathrm{d}t}$ can be calculated using a matrix-free iterative solver for system (2) or a truncated Neumann series approximation to $M_{C}^{-1}$. Such algorithms for inversion of mass matrices can be found in abgrall2017 ; donea1984 ; guermond2014 . For linear finite elements, which we are using in this work, it is worthwhile to approximate $\dot{u}_{i}$ by
$\dot{u}_{i}^{L}=\frac{1}{m_{i}}\sum_{e\in\mathcal{E}_{i}}[d^{e}(\bar{u}^{e}-u_{i})-\mathbf{f}(\bar{u}^{e})\cdot\mathbf{c}_{i}^{e}]=\frac{1}{m_{i}}\sum_{e\in\mathcal{E}_{i}}2d^{e}(\bar{u}_{i}^{e}-u_{i}).$
This approximation corresponds to adding a high-order stabilization term convex .
In the low-order method (3), the element contribution $r_{i}^{e,H}$ is replaced with
$r_{i}^{e,L}=r_{i}^{e,H}-f_{i}^{e},$
$f_{i}^{e}=\int_{K_{e}}\varphi_{i}(\dot{u}_{i}-\dot{u}_{h})\,\mathrm{d}\mathbf{x}+\int_{K_{e}}\nabla\varphi_{i}\cdot(\mathbf{f}(u_{h})-\mathbf{f}(\bar{u}^{e}))\,\mathrm{d}\mathbf{x}-d^{e}(\bar{u}^{e}-u_{i}).$
The Lagrange basis functions $\varphi_{i}$ have the property that $\sum_{i\in\mathcal{N}^{e}}\varphi_{i}\equiv 1$. Hence, $\sum_{i\in\mathcal{N}^{e}}\nabla\varphi_{i}\equiv\mathbf{0}$ and $\sum_{i\in\mathcal{N}^{e}}\int_{K_{e}}\varphi_{i}(\dot{u}_{i}-\dot{u}_{h})\,\mathrm{d}\mathbf{x}=0$. Furthermore, $\sum_{i\in\mathcal{N}^{e}}(\bar{u}^{e}-u_{i})=0$ by definition of $\bar{u}^{e}$. It follows that the element contributions satisfy
$\sum_{i\in\mathcal{N}^{e}}f_{i}^{e}=0,\qquad\sum_{i\in\mathcal{N}^{e}}r_{i}^{e,L}=\sum_{i\in\mathcal{N}^{e}}r_{i}^{e,H}=r^{e},$
$r^{e}=-\int_{K_{e}}\nabla\cdot\mathbf{f}(u_{h})\,\mathrm{d}\mathbf{x}=-\int_{\partial K_{e}}\mathbf{f}(u_{h})\cdot\mathbf{n}\,\mathrm{d}s,$ (6)
where $\mathbf{n}$ denotes the unit outward normal to $\partial K_{e}$. The general form (5) with
$r_{i}^{e}=r_{i}^{e,L}+f_{i}^{e,*}$
reduces to (2) for $f_{i}^{e,*}=f_{i}^{e}$ and to (3) for $f_{i}^{e,*}=0$. The fluctuation $r^{e}=\sum_{i\in\mathcal{N}^{e}}r_{i}^{e}$ defined in formula (6) is preserved provided that $\sum_{i\in\mathcal{N}^{e}}f_{i}^{e,*}=0$.
A classical RD method for a scalar conservation law of the form (1) constructs nodal residuals $r_{i}^{e}=\beta_{i}^{e}r^{e}$ using distribution weights $\beta_{i}^{e}$ such that $\sum_{i\in\mathcal{N}^{e}}\beta_{i}^{e}=1$. If $r^{e}\neq 0$ and $r_{i}^{e}$ is given, then $\beta_{i}^{e}=r_{i}^{e}/r^{e}$. For $r^{e}=0$, the assumption that $r_{i}^{e}=\beta_{i}^{e}r^{e}$ implies $r_{i}^{e}=0\ \forall i\in\mathcal{N}^{e}$. However, $r^{e}=0$ holds also if $r^{e}_{+}=-r^{e}_{-}\neq 0$, where
$r^{e}_{+}=\sum_{i\in\mathcal{N}^{e}}\max\{0,r_{i}^{e}\},\qquad r^{e}_{-}=\sum_{i\in\mathcal{N}^{e}}\min\{0,r_{i}^{e}\}.$
The authors of RD1 ; RD2 distribute the fluctuations $r^{e}_{\pm}$ using weights $\beta_{i,\pm}^{e}$ such that
$r_{i}^{e}=\begin{cases}\beta_{i,+}^{e}r^{e}_{+}&\mbox{if}\ r_{i}^{e}>0,\\ \beta_{i,-}^{e}r^{e}_{-}&\mbox{if}\ r_{i}^{e}<0,\end{cases}\qquad\beta_{i,\pm}^{e}\geq 0,\quad\sum_{i\in\mathcal{N}^{e}}\beta_{i,\pm}^{e}=1.$
As we show in the next section, such distribution can also be used for $(f_{i}^{e,*})_{i\in\mathcal{N}^{e}}$.
4 Convex limiting
Recall that the IDP property of the low-order scheme was shown by representing the result of the explicit update (4) as a convex combination of states belonging to a convex invariant set $\mathcal{G}$. Modern convex limiting techniques ensure the existence of such representations for high-order extensions. Element-based limiters of this kind can be subdivided into (A) flux-corrected transport (FCT) algorithms that add $f_{i}^{e,*}$ to a low-order predictor RD1 ; book ; lohmann2017 and (B) monolithic semi-discrete alternatives that insert $f_{i}^{e,*}$ into the right-hand side of the low-order scheme RD2 ; book ; preprint .
In the first stage of a type A scheme, a low-order IDP approximation $u_{h}^{L}$ is calculated using (4). The second stage corrects $u_{i}^{L}\in\mathcal{G}$ as follows:
$u_{i}^{\rm CL}=u_{i}^{L}+\frac{\Delta t}{m_{i}}\sum_{e\in\mathcal{E}_{i}}f_{i}^{e,*}=\frac{1}{m_{i}}\sum_{e\in\mathcal{E}_{i}}m_{i}^{e}\bar{u}_{i}^{e,*},\qquad\bar{u}_{i}^{e,*}=u_{i}^{L}+\frac{\Delta tf_{i}^{e,*}}{m_{i}^{e}}.$
Since $m_{i}=\sum_{e\in\mathcal{E}_{i}}m_{i}^{e}$ with $m_{i}^{e}>0$, the result $u_{i}^{\rm CL}$ is a convex combination of the auxiliary states $\bar{u}_{i}^{e,*}$. The element-based limiter should guarantee that $\bar{u}_{i}^{e,*}\in\mathcal{G}$. The first representative of such limiters was introduced in (cotter2016, , Sec. 2.3.2) in the context of flux-corrected remapping for scalar quantities. Element-based FCT/RD algorithms using this limiting strategy can be found in anderson2017 ; dobrev2018 ; RD1 ; kuzmin2020 ; lohmann2017 . An edge-based extension to nonlinear hyperbolic systems was proposed by Guermond et al. guermond2018 .
A monolithic convex limiting (MCL) algorithm is a single-stage type B scheme that avoids the computation of $u_{i}^{L}$ and replaces (4) with (cf. convex )
$u_{i}^{\rm CL}=\Big(1-\frac{2\Delta t}{m_{i}}\sum_{e\in\mathcal{E}_{i}}d^{e}\Big)u_{i}+\frac{2\Delta t}{m_{i}}\sum_{e\in\mathcal{E}_{i}}d^{e}\bar{u}_{i}^{e,*},\qquad\bar{u}_{i}^{e,*}=\bar{u}_{i}^{e}+\frac{f_{i}^{e,*}}{2d^{e}}.$
The IDP property is guaranteed under the CFL condition of the low-order scheme if $\bar{u}_{i}^{e,*}\in\mathcal{G}\ \forall e\in\mathcal{E}_{i}$. Representatives of element-based MCL schemes can be found in (book, , Chap. 4, 6). The original edge-based version was introduced in convex .
Limiters of type A and B have a lot in common. In both versions, $\bar{u}_{i}^{e,*}$ consists of a low-order IDP state $\bar{u}_{i}^{e,L}\in\{u_{i}^{L},\bar{u}_{i}^{e}\}$ and a correction term $f_{i}^{e,*}/\gamma_{i}^{e}$ with $\gamma_{i}^{e}\in\{m_{i}^{e}/\Delta t,2d^{e}\}$. Hence, the same algorithms can be used to enforce inequality constraints for $\bar{u}_{i}^{e,*}=\bar{u}_{i}^{e,L}+f_{i}^{e,*}/\gamma_{i}^{e}$ in type A and type B schemes convex ; book .
4.1 Limiting for scalars
In the scalar case, the states $\bar{u}_{i}^{e,*}$ can be constrained to stay in a local admissible range $u_{i}^{\min},u_{i}^{\max}]\subseteq[u^{\min},u^{\max}]=:\mathcal{G}$ such that $\bar{u}_{i}^{e,L}\in[u_{i}^{\min},u_{i}^{\max}]\ \forall e\in\mathcal{E}_{i}$. The constraints for element-based convex limiting of type A or B are formulated as follows:
$f_{i}^{e,\min}:=\gamma_{i}^{e}(u_{i}^{\min}-\bar{u}_{i}^{e,L})\leq f_{i}^{e,*}\leq\gamma_{i}^{e}(u_{i}^{\max}-\bar{u}_{i}^{e,L})=:f_{i}^{e,\max},$ (7)
$\sum_{i\in\mathcal{N}^{e}}f_{i}^{e,*}=0.$ (8)
The zero sum condition (8) is clearly satisfied for $f_{i}^{e,*}=\alpha^{e}f_{i}^{e}$, where $\alpha^{e}\in[0,1]$ is an element-based correction factor. Conditions (7) are satisfied for cotter2016
$\alpha^{e}=\min_{i\in\mathcal{N}^{e}}\alpha_{i}^{e},\qquad\alpha_{i}^{e}=\begin{cases}f_{i}^{e,\max}/f_{i}^{e}&\mbox{if}\ \ f_{i}^{e}>f_{i}^{e,\max},\\ f_{i}^{e,\min}/f_{i}^{e}&\mbox{if}\ \ f_{i}^{e}<f_{i}^{e,\min},\\ 1&\mbox{otherwise}.\end{cases}$ (9)
This scaling limiter was used in anderson2017 ; RD2 ; lohmann2017 to constrain finite element approximations of very high order. Interestingly enough, (9) has the same structure as the Barth–Jespersen slope limiter barth1989 for unstructured grid finite volume methods.
The application of $\alpha^{e}$ to $f_{i}^{e}=\beta_{i,\pm}^{e}f^{e}_{\pm}$ can be interpreted as residual distribution that scales the fluctuations $f^{e}_{\pm}$, while leaving the weights $\beta_{i,\pm}^{e}$ unchanged.
Another way to satisfy the limiting constraints (7) and (8) for each node $i\in\mathcal{N}^{e}$ is the following clip-and-scale (C&S) limiting strategy anderson2017 ; RD1 ; book ; lohmann2017 :
  1. 1.
    Calculate $\tilde{f}_{i}^{e}=\max\{f_{i}^{e,\min},\min\{f_{i}^{e},f_{i}^{e,\max}\}\}$ and the fluctuations
    $\tilde{f}^{e}_{+}=\sum_{i\in\mathcal{N}^{e}}\max\{0,\tilde{f}_{i}^{e}\},\qquad\tilde{f}^{e}_{-}=\sum_{i\in\mathcal{N}^{e}}\min\{0,\tilde{f}_{i}^{e}\}.$
  2. 2.
    Balance the sums of positive and negative components
    $f_{i}^{e,*}=\begin{cases}-(\tilde{f}^{e}_{-}/\tilde{f}^{e}_{+})\tilde{f}_{i}^{e}&\mbox{if}\ \tilde{f}_{i}^{e}>0,\ \tilde{f}^{e}_{+}+\tilde{f}^{e}_{-}>0,\\ -(\tilde{f}^{e}_{+}/\tilde{f}^{e}_{-})\tilde{f}_{i}^{e}&\mbox{if}\ \tilde{f}_{i}^{e}<0,\ \tilde{f}^{e}_{+}+\tilde{f}^{e}_{-}<0,\\ \tilde{f}_{i}^{e}&\mbox{otherwise}.\end{cases}$
The RD interpretation of this element-based limiter is as follows. The first step replaces $f_{i}^{e}=\beta_{i,\pm}^{e}f^{e}_{\pm}$ with clipped antidiffusive element contributions $\tilde{f}_{i}^{e}=\tilde{\beta}_{i,\pm}^{e}\tilde{f}^{e}_{\pm}$ that satisfy (7). The second step enforces the zero sum condition (8) by scaling the positive or negative fluctuation. The weights $\tilde{\beta}_{i,\pm}^{e}$ remain unchanged.
Note that $\tilde{f}_{i}^{e}=\tilde{\alpha}_{i}^{e}f_{i}^{e}$, where $\tilde{\alpha}_{i}^{e}$ corresponds to $\alpha_{i}^{e}$ defined by (9). Hence, the result of C&S limiting can be written as $f_{i}^{e,*}=\alpha_{i}^{e}f_{i}^{e}$, where
$\alpha_{i}^{e}=\begin{cases}-(\tilde{f}^{e}_{-}/\tilde{f}^{e}_{+})\tilde{\alpha}_{i}^{e}&\mbox{if}\ \tilde{f}_{i}^{e}>0,\ \tilde{f}^{e}_{+}+\tilde{f}^{e}_{-}>0,\\ -(\tilde{f}^{e}_{+}/\tilde{f}^{e}_{-})\tilde{\alpha}_{i}^{e}&\mbox{if}\ \tilde{f}_{i}^{e}<0,\ \tilde{f}^{e}_{+}+\tilde{f}^{e}_{-}<0,\\ \tilde{\alpha}_{i}^{e}&\text{otherwise}.\end{cases}$
Further examples of scalar element-based limiters can be found in (book, , Chap. 4).
4.2 Limiting for systems
Let us now consider a system of conservation laws for $u=(\varrho,\varrho\phi_{2},\ldots,\varrho\phi_{m})^{\top}$, where $\varrho=\rho\phi_{1}$ with $\phi_{1}=1$ is a density-like variable. The invariant set to be preserved by an IDP limiter for the element contributions $f_{i}^{e}=(f_{i,\varrho}^{e},f_{i,\varrho\phi_{1}}^{e},\ldots,f_{i,\varrho\phi_{m}}^{e})^{\top}$ is usually of the form $\mathcal{G}=\{u\in\mathbb{R}^{m}\,:\,\Phi_{1}(u)\geq 0,\ldots,\Phi_{M}(u)\geq 0\}$, where $\Phi_{1},\ldots,\Phi_{M}$ are quasi-concave scalar functions of the conserved variables.
In sequential convex limiting algorithms dobrev2018 ; book ; kuzmin2020 , local maximum principles for $\rho,\phi_{2},\ldots,\phi_{m}$ and the IDP property w.r.t. $\mathcal{G}$ are enforced as follows:
  1. 1.
    Given $f_{i,\varrho}^{e}$, calculate $f_{i,\varrho}^{e,*}$ using a scalar (scaling or C&S) limiter to ensure that
    $\varrho_{i}^{\min}\leq\bar{\varrho}_{i}^{e,*}=\bar{\varrho}_{i}^{e,L}+f_{i,\varrho}^{e,*}/\gamma_{i}^{e}\leq\varrho_{i}^{\max}.$
  2. 2.
    For $\phi\in\{\phi_{2},\ldots,\phi_{m}\}$, use a product rule version of the scalar limiter (see below) to construct element contributions $f_{i,\varrho\phi}^{e,*}\approx f_{i,\varrho\phi}^{e}$ satisfying
    $\bar{\varrho}_{i}^{e,*}\phi_{i}^{\min}\leq\bar{\varrho}_{i}^{e,*}\bar{\phi}_{i}^{e,*}=\overline{(\varrho\phi)}_{i}^{e,L}+f_{i,\varrho\phi}^{e,*}/\gamma_{i}^{e}\leq\bar{\varrho}_{i}^{e,*}\phi_{i}^{\max}.$
  3. 3.
    To keep $\bar{u}_{i}^{e,*}$ in the convex admissible set $\mathcal{G}$, apply $\alpha^{e}\in[0,1]$ such that
    $\bar{u}_{i}^{e}\in\mathcal{G}\quad\Rightarrow\quad\bar{u}_{i}^{e,*}=\bar{u}_{i}^{e,L}+\alpha^{e}f_{i}^{e,*}/\gamma_{i}^{e}\in\mathcal{G}.$
In Step 2, the local maximum principle $\phi_{i}^{\min}\leq\bar{\phi}_{i}^{e,*}\leq\phi_{i}^{\max}$ is enforced using a discrete version of $(\rho\phi)^{\prime}=\rho^{\prime}\phi+\rho\phi^{\prime}$. Simple scaling limiters designed for this purpose can be found in dobrev2018 ; kuzmin2020 . A product rule version of the C&S limiter was introduced in (book, , Sec. 6.3.2). It adjusts $f_{\varrho\phi}^{e}=(f_{i,\varrho\phi}^{e})_{i\in\mathcal{N}^{e}}$ in the following way:
  • Construct $\delta f_{\varrho\phi}^{e}=(\bar{\phi}_{i}^{e}f_{i,\varrho}^{e,*})_{i\in\mathcal{N}^{e}}$ using the nodal states $\bar{\phi}_{i}^{e}=\frac{\overline{(\varrho\phi)}_{i}^{e,L}}{\bar{\varrho}_{i}^{e,L}}$.
  • Apply the scaling operator $\mathcal{R}_{S}$ to $\delta f^{e}_{\varrho\phi}$ and calculate
    $g^{e}_{\varrho\phi}=f^{e}_{\varrho\phi}-\mathcal{R}_{S}(\delta f^{e}_{\varrho\phi}).$
  • Construct $\phi_{i}^{\min}=\min_{e\in\mathcal{E}_{i}}\bar{\phi}_{i}^{e,L}$ and $\phi_{i}^{\max}=\max_{e\in\mathcal{E}_{i}}\bar{\phi}_{i}^{e,L}$ using
    $\bar{\phi}_{i}^{e,L}=\frac{1}{\bar{\varrho}_{i}^{e,*}}\Big[\overline{(\varrho\phi)}_{i}^{e,L}+(f_{i,\varrho\phi}^{e}-g_{i,\varrho\phi}^{e})/\gamma_{i}^{e}\Big].$
  • Use the scalar C&S limiter to calculate $g_{i,\varrho\phi}^{e,*}$ such that
    $\bar{\varrho}_{i}^{e,*}\phi_{i}^{\min}\leq\bar{\rho}_{i}^{e,*}\bar{\phi}_{i}^{e,L}+g_{i,\varrho\phi}^{e,*}/\gamma_{i}^{e}\leq\bar{\varrho}_{i}^{e,*}\phi_{i}^{\max},\qquad\sum_{i\in\mathcal{N}^{e}}g_{i,\varrho\phi}^{e,*}=0.$
  • Calculate $f_{i,\varrho\phi}^{e,*}=f_{i,\varrho\phi}^{e}-g_{i,\varrho\phi}^{e}+g_{i,\varrho\phi}^{e,*}.$
If $\Phi_{1}=\varrho$ is the only variable that should stay nonnegative by definition of $\mathcal{G}$, then Step 3 of the sequential limiting procedure will use $\alpha^{e}=1$, because the local maximum principle $0\leq\varrho_{i}^{\min}\leq\bar{\varrho}_{i}^{e,*}\leq\varrho_{i}^{\max}$ is enforced in Step 1. Each additional variable $\Phi\in\{\Phi_{2},\ldots,\Phi_{M}\}$ imposes an upper bound $\alpha_{\Phi}^{e}$ on the value of $\alpha^{e}$. This bound can be calculated using a general line search algorithm guermond2018 or a closed-form expression derived from linear sufficient conditions convex ; book ; wu2023 . Limiters that ensure positivity preservation for the pressure and internal energy of the compressible Euler equations in this way can be found in convex ; kuzmin2020 ; abgrall2025 and (book, , Example 4.16).
In principle, it is possible to preserve local and/or global bounds for all scalar quantities of interest using the same scaling factor $\alpha^{e}$ for all components of $f_{i}^{e}$. However, such synchronized limiting is likely to produce more diffusive results.
4.3 Example: Euler equations
Figure 1 shows the low-order and MCL-C&S results for the double Mach reflection problem woodward . A detailed description of the computational setup can be found, e.g., in convex . The presented numerical solutions are nonoscillatory and satisfy the IDP constraints. The sequential MCL-C&S algorithm with IDP correction is clearly more accurate than the underlying element-based Rusanov method.
(a) low-order solution
(b) MCL-C&S solution
Figure 1: Double Mach reflection: density $\rho_{h}$ calculated using $\mathbb{P}_{1}$ elements and $h=1/128$.
5 Concluding remarks
The purpose of this review was to highlight the common structure of element-based convex limiting algorithms and their RD counterparts. Further examples of such FCT/MCL schemes can be found in (book, , Chapter 4). In our experience, recognition of existing relationships between seemingly different approaches leads to algorithms that advance the state of the art in all fields. For example, convex limiting techniques originally designed for discontinuous Galerkin (DG) methods can be adapted to the continuous finite element setting. In our recent work preprint , we propose an element-based MCL scheme that constrains intermediate cell averages and nodal states in much the same way as the Zhang–Shu limiter for the DG version. Moreover, we stabilize the antidiffusive element contributions using a dissipative component depending on a weighted essentially nonoscillatory (WENO) reconstruction. This nonlinear stabilization eliminates the need to preserve stringent local bounds. Hence, only the global bounds of IDP constraints need to be enforced using limiters.
Finally, we remark that the low-order scheme presented in Section 2 is equivalent to a subface finite volume (FV) discretization FV on a dual mesh $\mathcal{T}_{h}^{*}$ of control volumes associated with the nodal points $\mathbf{x}_{i}$ of a finite element mesh $\mathcal{T}_{h}$. In the FV version, vertices become cell centers, while cell centers become vertices (cf. selmin1993 ; selmin1996 ). This observation makes it possible to incorporate convex limiting of FCT and MCL type into FV schemes for general polygonal and polyhedral meshes. The virtual finite element MCL method proposed by Abgrall et al. pampa is also applicable to such meshes, and its derivation is based on the RD design philosophy.
Acknowledgements.
This work is dedicated to Professor Roland Duduchava (University of Georgia) in honor of his 80th anniversary. The author is grateful to him for the opportunity to give a talk on monolithic convex limiting at the Tbilisi Analysis & PDE online seminar in October 2025.